how much positive or negative charge is there in 1 litre of water
 

Grade:12

1 Answers

Deepak Kumar Shringi
5 years ago

moles in 1 lit of water=55.5
no of atoms = 55.5*Na where Na=avagadro no
in 1 atom no of proton=electron =10
charge on 1 electron =1.6* 10-19
total charge=1.6*10-19*55.5*6.023*1023*10
(positive or negative)=5.34*105